The reciprocal of 5 less than a # is 4 times the reciprocal of twice the #. find the #.
Thank you.

Answer by fractalier(2101)

Converting to algebra and calling the number x, we have
1 / (x - 5) = 4 / 2x
now cross-multiply and get
2x = 4(x - 5)
2x = 4x - 20
-2x = -20
x = 10